# Chemical Signatures of AGB Mass Transfer in Gaia White Dwarf Companions

### v1

@@ 8 Conclusions
  • We selected targets from the S. Shahaf et al. (2024) sample of systems classified as likely WD + MS binaries based on their Gaia astrometric orbits. We specifically selected several metal-poor targets with [Fe/H] <1.0<-1.0 (Section 2, Figure 1).
@@
  • We use ionization/excitation equilibrium of iron lines to determine Teff, log(g)(g), [Fe/H], and $\xi$. We find general agreement with existing literature values (Section 4.1, Figure 2).
@@
  • We measure abundances of iron (Fe I, Fe II), $\alpha$-elements (Mg I, Ca I, Si I), and s-process elements (Ba II, Y II) (Section 4.3, Figure 5). We identify 40 barium dwarfs ([Ba/Fe]>0.25>0.25\,dex OR [Y/Fe]>0.35>0.35\,dex), 36 of which are new discoveries (Section 5.1).
@@
  • About 25% of the likely WD+MS binaries in our sample are enriched in s-process elements. Their s-process abundances are uncorrelated with orbital period, in agreement with previous work. The incidence of s-process enhancement drops at high eccentricities and companion masses above $1.4\,M_{\odot}$, likely reflecting the presence of NS rather than WD companions. At lower companion masses, we find weak positive correlations with the component masses and an inverse correlation with the iron abundance (Section 5.2, Figure 8).
@@
  • We identify two barium dwarfs around companions whose masses may exceed the Chandrasekhar limit and which are in highly eccentric ($e>0.7$) orbits, making them possible MS + NS binaries. Further high-resolution spectroscopy with increased wavelength coverage containing additional s-process lines is necessary to strengthen constraints on their abundances, and multi-epoch radial velocities are needed to validate their orbits and confirm their nature (Section 5.3).
@@
  • We find evidence of strong C enhancement in three out of four of the most metal-poor barium dwarfs with $\mathrm{[Fe/H]}<-1.5$ (Figure 9), suggesting that there exists a smooth transition between barium dwarfs and CEMP-s stars (Section 5.4).
@@
  • Our sample roughly doubles the current number of barium dwarfs in AU-scale orbits. It extends the observed trend of [Ba/Fe] with [Fe/H] by P. Rekhi et al. (2025) to significantly more metal-poor stars with [Fe/H]<2.0<-2.0\,dex. (Section 6, Figure 10)
@@
  • Using theoretical AGB nucleosynthetic yields from the FRUITY database and 1D models of the companion MS stars to predict barium abundances in a simulated population of WD + MS binaries, we show that the observed diversity of s-process abundances can be attributed to the range in metallicities and component masses (which affect both the AGB yields and mass in the accretor’s convective envelope) and the age of the AGB donor at interaction. We also find that the fraction of stars that are barium dwarfs is sensitive to the assumed orbital evolution via MT. In general, systems for which MT began at initially wide separations with evolved AGB donors end up with greater barium enhancements (Section 7, Figure 14).
@@
Future data releases from the Gaia mission will build on this sample, extending to longer orbital periods and allowing for a more homogeneous selection of these systems to conduct more complete and detailed population-level studies.
@@
We thank Ana Escorza, Sophie van Eck, Alain Jorissen, and Param Rekhi for useful conversations. We are grateful to Yuri Beletsky, Sam Kim, Angela Hempel, and Maren Hempel for observing help. This research benefited from discussions in the ZTF Theory Network, funded in part by the Gordon and Betty Moore Foundation through Grant GBMF5076. NY acknowledges support from the Ezoe Memorial Recruit Foundation scholarship. This research was supported by NSF grant AST-2307232. {contribution}NY was responsible for completing the data analysis and population modeling of the systems in this work as well as writing and submitting the manuscript. KE advised NY throughout the project, from initial conceptualization to editing the manuscript. HR provided expertise on the spectral analysis and ensured that the stellar parameters and abundances obtained were reliable. RA re-analyzed Gaia XP spectra using corrected parallaxes appropriate for our systems, which we used in our analysis. $S$ created the original sample of astrometric binaries from which our systems were selected.
@@ Appendix A Appendix information
A.1 Results of stellar parameter determination
@@
In Figure 15, we show examples of the final output from q2 in measuring stellar parameters via excitation/ionization equilibrium for three targets with varying degrees of success (Section 4.1). Successful convergence is achieved when there are minimal trends in the implied abundances of both Fe I and Fe II as a function of several line properties, plotted in each of the three rows. Having a small number of lines with reliably measured EWs as a result of poor SNR and/or rotational broadening can lead to convergence failure. For a typical spectrum of our sample, we find that our method is effective for $v\sin(i)\lesssim 20\,$km/s. Detailed descriptions for the three examples can be found in the figure caption.
@@
Figure 15: Results of attempting to achieve excitation/ionization equilibrium using iron lines with q2 (Section 4.1). The blue cross markers correspond to Fe I, while the green circle markers correspond to Fe II lines. The best fit linear model is also plotted in each panel. We show an example of a spectrum which (A) achieves convergence with reliable results on the first attempt, (B) requires fixing log(g)(g) to its initial value to avoid it from going to $5.0$, and (C) fails to achieve convergence due to rotationally broadened lines. Note that the scatter in the data depends on the SNR and resolution of each spectrum.
A.2 Effect of systematic uncertainty on $T_{\rm eff}$
@@
As discussed in Section 4.1, the best-fit $T_{\rm eff}$ values derived from high-resolution spectra deviate from those of R. Andrae et al. (2023) which are measured using Gaia XP spectra and broad-band photometry. In most cases, the SEDs predicted using $T_{\rm eff}$ values from the latter are more consistent with photometric observations. Nevertheless, we use the spectroscopic stellar parameters to derive our fiducial abundances. Here, we re-derive them using $T_{\rm eff}$ values from R. Andrae et al. (2023), and in Figure 16, we compare the resulting [Ba/H] (as opposed to [Ba/Fe] which also depends on the iron abundance). We find that our measurements are generally higher, and that this discrepancy increases for cooler stars. However, the discrepancy is less than 0.30.3\,dex for the bulk of the sample. Using $T_{\rm eff}$ values from R. Andrae et al. (2023) would only reduce the total number of stars classified to be barium dwarfs by one (Gaia DR3 source ID 3474329544020730496, with a [Ba/Fe] close to the lower bound of 0.25 dex).
@@
Figure 16: Left: Comparison of barium abundances derived taking the best-fit $T_{\rm eff}$ from this work vs. the value from R. Andrae et al. (2023) (modified to use parallaxes corrected for binary motion; Section 4.1). We take our best-fit values for all other stellar parameters ([Fe/H], log(g)(g), and $\xi$). Right: The difference in the abundances as a function of $T_{\rm eff}$. The difference increases for cooler stars, where the two $T_{\rm eff}$ values become increasingly discrepant.
A.3 Tables of atmospheric parameters and abundances
@@
In Tables 2 and 3, we provide snippets of tables summarizing the derived atmospheric parameters and abundances. The full tables containing information for all 160 stars in our final sample can be downloaded from the online journal.
@@
Table 2: Atmospheric parameters for systems which achieved successful convergence (Section 4.1). The complete table can be downloaded online.
@@
Table 3: Absolute abundances derived for systems in Table 2 (Section 4.3). The reported errors are the standard deviations of abundances inferred from multiple lines divided by square root of the number of lines (NN). In the case of a single line, we adopt a fixed error of 0.2 dex. The complete table can be downloaded online.
A.4 Trends with yttrium abundance
Figure 17 plots [Y/Fe] as a function of several orbital and stellar parameters, analogous to that of [Ba/Fe] in Figure 8 where similar trends (or lack thereof) are found (Section 5.2).

### v2

@@ VIII Conclusions
  • We selected targets from the Shahaf et al. (2024) sample of systems classified as likely WD + MS binaries based on their Gaia astrometric orbits. We specifically selected several metal-poor targets with [Fe/H] <1.0<-1.0 (Section II, Figure 1).
@@
  • We use ionization/excitation equilibrium of iron lines to determine Teff, log(g)(g), [Fe/H], and $\xi$. We find general agreement with existing literature values (Section IV.1, Figure 2).
@@
  • We measure abundances of iron (Fe I, Fe II), $\alpha$-elements (Mg I, Ca I, Si I), and s-process elements (Ba II, Y II) (Section IV.3, Figure 5). We identify 43 barium dwarfs ([Ba/Fe]>0.25>0.25\,dex OR [Y/Fe]>0.35>0.35\,dex), 39 of which are new discoveries (Section V.1).
@@
  • About 25% of the likely WD+MS binaries in our sample are enriched in s-process elements. Their s-process abundances are uncorrelated with orbital period, in agreement with previous work. The incidence of s-process enhancement drops at high eccentricities and companion masses above $1.4\,M_{\odot}$, likely reflecting the presence of NS rather than WD companions. At lower companion masses, we find weak positive correlations with the component masses and an inverse correlation with the iron abundance (Section V.2, Figure 8).
@@
  • We identify two barium dwarfs around companions whose masses may exceed the Chandrasekhar limit and which are in highly eccentric ($e>0.7$) orbits, making them possible MS + NS binaries. Further high-resolution spectroscopy with increased wavelength coverage containing additional s-process lines is necessary to strengthen constraints on their abundances, and multi-epoch radial velocities are needed to validate their orbits and confirm their nature (Section V.3).
@@
  • We find evidence of strong C enhancement in three out of four of the most metal-poor barium dwarfs with $\mathrm{[Fe/H]}<-1.5$ (Figure 9), suggesting that there exists a smooth transition between barium dwarfs and CEMP-s stars (Section V.4).
@@
  • Our sample roughly doubles the current number of barium dwarfs in AU-scale orbits. It extends the observed trend of [Ba/Fe] with [Fe/H] by Rekhi et al. (2025) to significantly more metal-poor stars with [Fe/H]<2.0<-2.0\,dex. (Section VI, Figure 10)
@@
  • Using theoretical AGB nucleosynthetic yields from the FRUITY database and 1D models of the companion MS stars to predict barium abundances in a simulated population of WD + MS binaries, we show that the observed diversity of s-process abundances can be attributed to the range in metallicities and component masses (which affect both the AGB yields and mass in the accretor’s convective envelope) and the age of the AGB donor at interaction. We also find that the fraction of stars that are barium dwarfs is sensitive to the assumed orbital evolution via MT. In general, systems for which MT began at initially wide separations with evolved AGB donors end up with greater barium enhancements (Section VII, Figure 14).
@@
Further analysis to determine precise C abundances as well as additional spectra covering a wider range of wavelengths and elemental lines would be beneficial follow-up work. Moreover, future data releases from the Gaia mission will build on this sample, extending to longer orbital periods and allowing for a more homogeneous selection of these systems to conduct more complete and detailed population-level studies.
@@
We thank the anonymous referee for detailed feedback which improved the manuscript. We thank Ana Escorza, Sophie van Eck, Alain Jorissen, Param Rekhi, Ben Roulston, and Jay Farihi for useful conversations. We are grateful to Yuri Beletsky, Sam Kim, Angela Hempel, and Maren Hempel for observing help. This research benefited from discussions in the ZTF Theory Network, funded in part by the Gordon and Betty Moore Foundation through Grant GBMF5076. NY acknowledges support from the Ezoe Memorial Recruit Foundation scholarship. This research was supported by NSF grant AST-2307232.
{contribution}
NY was responsible for completing the data analysis and population modeling of the systems in this work as well as writing and submitting the manuscript. KE advised NY throughout the project, from initial conceptualization to editing the manuscript. HR provided expertise on the spectral analysis and ensured that the stellar parameters and abundances obtained were reliable. RA re-analyzed Gaia XP spectra using corrected parallaxes appropriate for our systems, which we used in our analysis. SS created the original sample of astrometric binaries from which our systems were selected.
@@ Appendix A Appendix information
A.1 Radial velocities
We summarize RVs measured for each of our spectra in Table 2. Errors are calculated by iSpec following Zucker (2003). The full table can be found in the supplementary materials.
We caution that we do not attempt to account for zero-point offsets between different instruments, which are expected to be of order $1\,{\rm km\,s^{-1}}$. Precision is also limited by the fact that we used a fixed template at solar metallicity for the cross-correlation across all spectra.
Gaia DR3 source ID JD RV Instrument
41408333753757056 2460161.0915 -10.6 $\pm$ 2.1 HIRESr
220012968211559296 2460161.0787 -24.5 $\pm$ 2.3 HIRESr
352523017213180544 2460879.0566 20.0 $\pm$ 2.8 HIRESb
369419178036177792 2460879.1283 -65.5 $\pm$ 2.5 HIRESb
383389126103078400 2460847.0807 -148.4 $\pm$ 3.7 HIRESb
Table 2: Measured RVs for all spectra. The complete table can be downloaded online.
A.2 Stellar parameters from Gaia XP spectra
In Table 3, we provide stellar parameters re-calculated using methods from Andrae et al. (2023) but with parallaxes corrected for binary motion. We used these values as initial guesses in our analysis (Section IV.1).
Gaia DR3 source ID $T_{\rm eff}$ log($g$) [Fe/H]
2808994137268673152 6255 $\pm$ 80 4.29 $\pm$ 0.03 -0.15 $\pm$ 0.08
2124249013097682944 6211 $\pm$ 97 4.26 $\pm$ 0.04 -1.88 $\pm$ 0.16
2097059980331156480 5594 $\pm$ 54 4.50 $\pm$ 0.01 -0.34 $\pm$ 0.05
2795884385252655232 6027 $\pm$ 90 4.33 $\pm$ 0.03 -0.63 $\pm$ 0.08
2105234402606388352 5400 $\pm$ 42 4.53 $\pm$ 0.01 -0.47 $\pm$ 0.06
Table 3: Stellar parameters derived from Gaia XP spectra re-calculated using methods from Andrae et al. (2023) with parallaxes from the nss_two_body_orbit table. The complete table can be downloaded online.
A.3 Results of stellar parameter determination
@@
In Figure 15, we show examples of the final output from q2 in measuring stellar parameters via excitation/ionization equilibrium for three targets with varying degrees of success (Section IV.1). Successful convergence is achieved when there are minimal trends in the implied abundances of both Fe I and Fe II as a function of several line properties, plotted in each of the three rows. Having a small number of lines with reliably measured EWs as a result of poor SNR and/or rotational broadening can lead to convergence failure. For a typical spectrum of our sample, we find that our method is effective for $v\sin(i)\lesssim 20\,$km/s. Detailed descriptions for the three examples can be found in the figure caption.
@@
Figure 15: Results of attempting to achieve excitation/ionization equilibrium using iron lines with q2 (Section IV.1). The blue cross markers correspond to Fe I, while the green circle markers correspond to Fe II lines. The best fit linear model is also plotted in each panel. We show an example of a spectrum which (A) achieves convergence with reliable results on the first attempt, (B) requires fixing log(g)(g) to its initial value to avoid it from going to $5.0$, and (C) fails to achieve convergence due to rotationally broadened lines. Note that the scatter in the data depends on the SNR and resolution of each spectrum.
A.4 Effect of systematic uncertainty on $T_{\rm eff}$
@@
As discussed in Section IV.1, the best-fit $T_{\rm eff}$ values derived from high-resolution spectra deviate from those of Andrae et al. (2023) which are measured using Gaia XP spectra and broad-band photometry. In most cases, the SEDs predicted using $T_{\rm eff}$ values from the latter are more consistent with photometric observations. Nevertheless, we use the spectroscopic stellar parameters to derive our fiducial abundances. Here, we re-derive them using $T_{\rm eff}$ values from Andrae et al. (2023), and in Figure 16, we compare the resulting [Ba/H] (as opposed to [Ba/Fe] which also depends on the iron abundance). We find that our measurements are generally higher, and that this discrepancy increases for cooler stars. However, the discrepancy is less than 0.30.3\,dex for the bulk of the sample. Using $T_{\rm eff}$ values from Andrae et al. (2023) would only reduce the total number of stars classified to be barium dwarfs by one (Gaia DR3 source ID 3474329544020730496, with a [Ba/Fe] close to the lower bound of 0.25 dex).
@@
Figure 16: Left: Comparison of barium abundances derived taking the best-fit $T_{\rm eff}$ from this work vs. the value from Andrae et al. (2023) (modified to use parallaxes corrected for binary motion; Section IV.1). We take our best-fit values for all other stellar parameters ([Fe/H], log(g)(g), and $\xi$). Right: The difference in the abundances as a function of $T_{\rm eff}$. The difference increases for cooler stars, where the two $T_{\rm eff}$ values become increasingly discrepant.
A.5 Tables of atmospheric parameters and abundances
@@
In Tables 4 and 5, we provide snippets of tables summarizing the derived atmospheric parameters and abundances. The full tables containing information for all 160 stars in our final sample can be downloaded from the online journal.
@@
Table 4: Atmospheric parameters for systems which achieved successful convergence (Section IV.1). The complete table can be downloaded online.
@@
Table 5: Absolute abundances derived for systems in Table 4 (Section IV.3). The reported errors are the standard deviations of abundances inferred from multiple lines divided by square root of the number of lines (NN). In the case of a single line, we adopt a fixed error of 0.2 dex. The complete table can be downloaded online.
A.6 Trends with yttrium abundance
Figure 17 plots [Y/Fe] as a function of several orbital and stellar parameters, analogous to that of [Ba/Fe] in Figure 8 where similar trends (or lack thereof) are found (Section V.2).